OpenStudy (anonymous):

please help!! solve the integral of (cos^2)(sin^4)

OpenStudy (anonymous):

put cos^2=c0s2x+1

OpenStudy (anonymous):

or c0s^2=1-sin^@

OpenStudy (anonymous):

but then how do you solve the integral of sin^4??

OpenStudy (anonymous):

first expand sin^4(x) in terms of sin^2 and cos^2

OpenStudy (anonymous):

@chenna, how to expand it??

OpenStudy (anonymous):

put sin^4=sin^2sin^2 then=sin^(1-cos^)

OpenStudy (anonymous):

sry sin^2(1-cos^2)

OpenStudy (anonymous):

vipul. you are wrong. http://www.stewartcalculus.com/data/CALCULUS%20Concepts%20and%20Contexts/upfiles/3c3-TrigonometIntegrals_Stu.pdf follow this link. this will help you.

OpenStudy (anonymous):

if you can't follow ask me. message me.

OpenStudy (anonymous):

@chenna well now i got the integral of sin^4....but can u help me solve the integral of sin^6??

OpenStudy (anonymous):

first you write sin^4(x) and cos^2(x) interms of cos2x

OpenStudy (anonymous):

make use of example 4 in that link i provided you.

OpenStudy (anonymous):

first write cos^2(x) in terms of cos2x. what do you get?

OpenStudy (anonymous):

any progress? ask me if you don't get it.